OpenStudy (anonymous):

solve the equation by making an appropriate substitution? 2x-9 x(sq rt) -5=0 How do you do this?

OpenStudy (anonymous):

\[2x - 9\sqrt{x} -5 = 0\] Is it right?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

I take that back it is -2x not 2x, sorry..

OpenStudy (anonymous):

My bad, you were right the first time. It' time to go to bed, I'm getting a brain-freeze... it is 2x

OpenStudy (anonymous):

The substitution you want to make is: \[y = \sqrt{x} \Rightarrow y^{2} = x\] that way your equation becomes: \[2x-9\sqrt{x}-5 =0\Rightarrow 2y^{2}-9y-5 = 0\] Then you would use the quadratic formula

OpenStudy (anonymous):

if someone else could finish, it would be greatly appreciated lol, im in the middle of eating and watching Sucker Punch >.>

myininaya (myininaya):

2(-5)=-10(1) -10+1=-9 we will replace -9y with -10y+1y \[2y^2-10y+1y-5=0\] \[2y(y-5)+1(y-5)=0\] \[(y-5)(2y+1)=0\] \[y=5, y=\frac{-1}{2}\] but remember \[y=\sqrt{x}\] so we have \[\sqrt{x}=5, \sqrt{x}=\frac{-1}{2}\] to solve this we square both sides \[x=25,x=\frac{1}{4}\] we should check these since we raised both sides to an even power

myininaya (myininaya):

so lets check x=25 \[2*25-9\sqrt{25}-5=50-9*5-5=50-45-5=5-5=0\] since we have the samething on both sides then x=25 is a solution and lets check x=1/4 \[2*\frac{1}{4}-9\sqrt{\frac{1}{4}}-5=\frac{1}{2}-9*\frac{1}{2}-5=\frac{-8}{2}-5=-4-5=-9\] and x=1/4 is not a solution since both sides are not the same